# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ 3.1 Main Theorem
One important implication of our result is that it addresses an open problem about NC. Specifically, almost all existing works assume the unconstrained feature model [84, 91] (or layer-peeled model [38, 61]) to analyze the NC phenomenon, where the last-layer features of the deep network are treated as free optimization variables to simplify the interactions across layers; see, e.g., [127, 61, 116, 130, 125, 128, 80]. However, a major drawback of this model is that it overlooks the hierarchical structures of deep networks as well as the structure of the input data. In this work, we address this issue without assuming the unconstrained feature model. Specifically, according to 1, when a linear network is deep enough and trained on nearly orthogonal data, the last-layer features within the same class concentrate around their class means, while the last-layer features from different classes are nearly orthogonal to each other. This directly implies the last-layer features exhibit variability collapse and convergence to simplex equiangular right frame approximately after centralization.
Since the submission of our manuscript, several follow-up works have investigated NC beyond the unconstrained feature model in different settings; see, e.g., [49, 115, 62, 69, 119]. In particular, [49] studied 2-layer and 3-layer ReLU neural networks and investigated the impact of network width, depth, data dimension, and statistical properties of input data on NC. [115] characterized the geometric properties of intermediate layers of ResNet and proposed a conjecture on progressive feedforward collapse. Recently, [62] studied a hybrid network with nonlinear layers followed by at least two linear layers and provided generic guarantees on progressive NC. It is worth noting that their derived conditions extend those in 2 and are slightly more general.
(a) A 7-layer MLP network with ReLU
(b) A 7-layer hybrid network
Figure 5: Within-class compression and between-class discrimination of features of two 7-layer networks. We use the same setup as in Figure 1 and plot the metrics of within-class compression and between-class discrimination (see Definition 1) of the two networks across layers after training, respectively. More experimental details are deferred to Section 5.1.1.
@@
Progressive feature compression and discrimination in 1 provide guiding principles for network architecture design. Specfically, according to (14), (15), and (16), the features are more compressed within the same class and more discriminated between different classes, improving the separability of input data as the depth of networks increases. This is also supported by our experiments in Figure 12. This indicates that the networks should be deep enough for effective data separation in classification problems. However, it is worth noting that the belief that deeper networks are better is not always true. Indeed, it becomes increasingly challenging to train a neural network as it gets deeper, especially for DLNs [41]. Moreover, over-compressed features may degrade the out-of-distribution performance of deep networks as shown in [83]. This, together with our result, indicates that neural networks should not be too deep for improved out-of-distribution generalization performance.
@@
In contrastive learning [21, 20], a successful empirical approach to improving transfer learning performance in downstream tasks involves the use of projection heads [79]. These projection heads, typically consisting of one or several MLP layers, are added between the feature extractor and the final classifier layers during pre-training. For downstream tasks, the projection head is discarded and only the features learned by the feature extractor are utilized for transfer learning. Recent works [79, 66, 44, 118] have established an empirical correlation between the degree of feature collapse during pre-training and downstream performance: less feature collapse leads to more transferable models. However, it remains unclear why features prior to projection heads exhibit less collapse and greater transferability. Our study addresses this question and offers a theoretical insight into the utilization of projection heads. According to 1, it becomes apparent that features from the final layers tend to be more collapsed than the features before the projection head. This, together with the empirical correlation between feature collapse and transferability, implies that leveraging features before projection heads improves transferability. This is also supported by our experimental results in Section 5.3.2. Moreover, the progressive compression pattern in 1 also provides insight into the phenomenon studied in [122], which suggests that deeper layers in a neural network become excessively specialized for the pre-training task, consequently limiting their effectiveness in transfer learning.
3.2 Discussions on 2
In this subsection, we justify the properties imposed on the trained network weights as outlined in 2, drawing insights from both our theoretical and experimental findings. It is important to emphasize that the solution described in 2 represents just one of infinitely many globally optimal solutions for Problem (3). Nevertheless, because of the implicit bias in GD training dynamics, GD iterations with proper initialization almost always converge to the desired solutions outlined in 2. In the subsequent discussion, we will delve into this phenomenon in greater depth.
@@ 3.2 Discussions on 2
[…]
where $\bm{U}\in\mathcal{O}^{K}$. It is worth noting that orthogonal initialization is widely used to train deep networks, which can speed up the convergence of GD; see, e.g., PSG [93], SMG [99], XBSD+ [117], HXP [55].
Theoretical justification of 2.
Theoretically, we can prove the conditions outlined in Assumption 2 when the gradient flow is trained on a square and orthogonal data matrix using the results in [3, 124].
Proposition 1.
@@
The proof of this proposition is deferred to Section B.5. Although we can only prove 2 in this restrictive setting, it actually holds in more general settings. We substantiate this claim by drawing support from empirical evidence and existing findings in the literature, as elaborated in the following.
Empirical justifications of 2.
@@
Figure 6: 2 holds approximately among different network configurations. We train DLNs using the orthogonal initialization with varying widths and depths. For each network width $d$ and depth $L$, we show the minimum-norm residual, the average weight balancedness residual of all weights, and the variance of the singular values of all weights (see (5.2.1) for details) for the DLNs that have been trained until convergence. Each data point represents the mean value of 10 runs with different random seeds. Here, we plot these metrics in (a) with fixed network width $d$ and varying depth $L$ and (b) with fixed network depth $L=5$ and varying network width $d$. More experimental details are deferred to Section 5.2.1.
Prior arts on implicit bias support 2.
Moreover, 2 is also well supported by many related results in the literature.
  • Minimum-norm solutions and balanced weights. It has been extensively studied in the literature that GD with proper initialization converges to minimum-norm and balanced solutions, especially in the setting of gradient flow [3, 22, 31, 86]. Specifically, [31, 3] proved that for linear networks, the iterates of gradient flow satisfy
@@
    where $t$ denotes a continuous time index. This, together with (18), implies that (11) holds with $\delta=\xi\sqrt{d-K}$. Moreover, using the result in [86], if we initialize the weights of two-layer networks as in (18) with $\bm{W}_{2}(0)\bm{W}_{1}(0)\in\mathrm{span}(\bm{X})$, gradient flow always yields the minimum-norm solution upon convergence. However, the conditions (10) and (11) are rarely studied in the context of GD for training deep networks. In this context, we empirically verify these two conditions as shown in Figure 6 under different network configurations.
  • Approximate low-rankness of the weights. Recently, numerous studies have demonstrated that GD exhibits a bias towards low-rank weights [4, 46, 124, 72]. In particular, [124, 72] showed that when the input data $\bm{X}$ is orthonormal, learning dynamics of GD for DLNs in (17), with the initialization in (18), only updates an invariant subspace of dimension $2K$ for each weight matrix across all layers, where the invariant subspace is spanned by the singular vectors associated with the $K$-largest and $K$-smallest singular values, and the rest singular values remain unchanged. More experimental demonstration can be found in Appendix C.
@@ 4 Relationship to Prior Arts
Deep networks, organized in hierarchical layers, can perform effective and automatic feature learning [10], where the layers learn useful representations of the data. To better understand hierarchical feature learning, plenty of studies have been conducted to investigate the structures of features learned at intermediate layers. One line of these works is to investigate the neural collapse (NC) properties at intermediate layers. For example, [110] extended the study of neural collapse to three-layer nonlinear networks with the MSE loss, showing that the features of each layer exhibit neural collapse. [33] generalized the study of neural collapse for DLNs with imbalanced training data, and they draw a similar conclusion as [110] that the features of each layer are collapsed. However, these results are based on the unconstrained feature model and thus cannot capture the input-output relationship of the neural networks. Moreover, the conclusion they draw on the collapse of intermediate features is far from what we observe in practice. Indeed, it is observed on both linear and nonlinear networks in Figure 4 that the intermediate features are progressively compressed across layers rather than exactly collapsing to their means for each layer. In comparison, under the assumption that the input $\bm{X}$ is nearly orthogonal in 1, our result characterizes the progressive compression and discrimination phenomenon to capture the training phenomena on practical networks as demonstrated in Figure 4.
On the other hand, another line of work [123, 24, 27, 85, 121] argues that deep networks prevent within-class feature compression while promoting between-class feature discrimination, contrasting with our study where features are compressed across layers. Such a difference comes from the choices of loss functions. In our work, we focused on the study of the commonly used MSE loss, with the goal of understanding a prevalent phenomenon in classical training of deep networks. It has been empirically shown that increasing feature compression is beneficial for improving in distribution generalization and robustness [91, 79, 17, 53, 19]. In comparison, [123] introduced a new maximum coding rate reduction loss that is intentionally designed to prevent feature compression.
Distinction from the neural tangent kernel: feature learning.
There has been a rich line of work that studies mathematical insights into deep networks in the regime where the width of a network is sufficiently large. In the infinite-width limit, it has been demonstrated that neural networks converge to a solution to a kernel least squares problem [60]. This is referred to as the neural tangent kernel (NTK), which serves as an important theoretical tool in understanding the convergence [11, 30, 34] and generalization [8, 5] of neural networks in the kernel regime. However, its effectiveness is limited when it comes to understanding hierarchical structure and justifying the advantages of deep architectures in neural networks. For example, recent works [9, 18, 32, 82] have identified sets of functions that deep architectures can efficiently learn, yet kernel methods fall short in learning these functions in theory. In comparison, our work offers insights into understanding hierarchical representations of neural networks in terms of feature compression and discrimination from shallow to deep layers.
@@
Our main result in 1 is based on 2 for the trained weights. As discussed in Section 3.2, 2 holds as a consequence of results in recent works on analyzing the learning dynamics and implicit bias of gradient flow or GD in training deep networks. We briefly review the related results as follows. For training DLNs, [2] established linear convergence of GD based upon whitened data and with a similar setup to ours. For Gaussian initialization, [30] showed that GD also converges globally at a linear rate when the width of hidden layers is larger than the depth, whereas [55] demonstrated the advantage of orthogonal initialization over random initialization by showing that linear convergence of GD with orthogonal initialization is independent of the depth. More recent developments can be found in [42, 90, 97] for studying GD dynamics. On the other hand, another line of works focused on studying gradient flow for learning DLNs due to its simplicity [16, 36, 86, 111, 16], by analyzing its convergence behavior.
Numerous studies have shown that the effectiveness of deep learning is partially due to the implicit bias of its learning dynamics, which favors some particular solutions that generalize exceptionally well without overfitting in the over-parameterized setting [13, 51, 87]. To gain insight into the implicit bias of GD for training deep networks, a line of recent work has shown that GD tends to learn simple functions [26, 46, 63, 71, 105, 113]. For instance, some studies have shown that GD is biased towards max-margin solutions in linear networks trained for binary classification via separable data [98, 71]. In addition to the simplicity bias, another line of work showed that deep networks trained by GD exhibit a bias towards low-rank solutions [46, 51, 124]. The works [42, 4] demonstrated that adding depth to a matrix factorization enhances an implicit tendency towards low-rank solutions, leading to more accurate recovery.
@@ 5 Experimental Results
In this section, we conduct various numerical experiments to validate our assumptions, verify the theoretical results, and investigate the implications of our results on both synthetic and real data sets. All of our experiments are conducted on a PC with 8GB memory and Intel(R) Core i5 1.4GHz CPU, except for those involving large datasets such as CIFAR and FashionMNIST, which are conducted on a computer server equipped with NVIDIA A40 GPUs. Our codes are implemented in Python and are available at https://github.com/Heimine/PNC_DLN. Throughout this section, we will repeatedly use multilayer perception (MLP) networks, where each layer of MLP networks consists of a linear layer and a batch norm layer [59] followed by ReLU activation.
The remaining sections are organized as follows: In Section 5.1, we provide detailed experimental setups for the results discussed in Section 1 and shown in Figures 5, 7, and 8. In Section 5.2, we provide experimental results to support 1 in Section 3 and validate Assumption 2 discussed in Section 3.2. Finally, we empirically explore the implications of our results in Section 5.3 beyond our assumptions.
@@ 5.1 Significance of Studying DLNs in Feature Learning
(a) A 7-layer MLP network with ReLU
(b) A 7-layer hybrid network
Figure 7: Accuracy and numerical rank of two 7-layer networks before training. We use the same setup as in Figures 1 and 5 but instead plot the numerical rank and training accuracy of the two networks across layers before training.
(a) A 7-layer MLP network with ReLU
(b) A 7-layer hybrid network
Figure 8: Within-class compression and Between-class discrimination of two 7-layer networks before training. We use the same setup as in Figure 1 and Figure 5 but instead plot the compression and discrimination statistics of the two networks across layers before training.
In this subsection, our goal is to empirically demonstrate the significance of studying DLNs in feature learning as discussed in Section 1. To achieve this, we investigate (i) the roles of linear layers and MLP layers at deep layers in feature learning and (ii) the importance of the depth of DLNs for generalization performance.

### v4

@@ 3.1 Main Theorem
One important implication of our result is that it addresses an open problem about NC. Specifically, almost all existing works assume the unconstrained features model (Mixon2020; papyan2020prevalence) to analyze the NC phenomenon, where the last-layer features of the deep network are treated as free optimization variables to simplify interactions across layers; see, e.g., zhu2021geometric; wang2022linear; zhou2022all; yaras2022neural; zhou2022optimization; li2023neural. However, a major drawback of this model is that it overlooks the hierarchical structure of deep networks as well as the structure of the input data. In this work, we address this issue without assuming the unconstrained feature model. Specifically, according to Theorem 4, for a sufficiently deep linear network trained on nearly orthogonal data, the last-layer features within the same class concentrate around their class means, while the last-layer features from different classes are nearly orthogonal to each other. This directly implies that the last-layer features exhibit variability collapse and convergence to simplex equiangular right frame approximately after centralization.
Since the submission of our manuscript, several follow-up works have investigated NC beyond the unconstrained feature model in different settings; see, e.g., hong2024beyond; wang2024progressive; jacot2025wide; kothapalli2024kernel; xu2023dynamics. In particular, hong2024beyond studied 2-layer and 3-layer ReLU neural networks and investigated the impact of network width, depth, data dimension, and statistical properties of input data on NC. wang2024progressive characterized the geometric properties of intermediate layers of ResNet and proposed a conjecture on progressive feedforward collapse. Recently, jacot2025wide studied a hybrid network with nonlinear layers followed by at least two linear layers and provided generic guarantees on progressive NC. It is worth noting that their derived conditions extend those in Assumption 3 and are slightly more general.
@@
Progressive feature compression and discrimination in Theorem 4 provide guiding principles for network architecture design. Specifically, according to (14), (15), and (16), features are more compressed within the same class and more discriminated between different classes, improving the separability of input data as the depth of the network increases. This is also supported by our experiments in Figure 4. This indicates that the network should be deep enough for effective data separation in classification problems. However, it is worth noting that the belief that deeper networks are better is not always true. Indeed, it becomes increasingly challenging to train a neural network as it gets deeper, especially for DLNs (glorot2010understanding). Moreover, over-compressed features may degrade the out-of-distribution performance of deep networks as shown in masarczyk2023tunnel. This, together with our result, indicates that neural networks should not be too deep for improved out-of-distribution generalization performance.
@@
In contrastive learning (chen2020simple; chen2021exploring), a successful empirical approach to improving transfer learning performance in downstream tasks involves the use of projection heads (li2022principled). These projection heads, typically consisting of one or several MLP layers, are added between the feature extractor and the final classifier layers during pre-training. For downstream tasks, the projection head is discarded and only the features learned by the feature extractor are utilized for transfer learning. Recent works (li2022principled; kornblith2021why; galanti2022on; xie2022hidden) have established an empirical correlation between the degree of feature collapse during pre-training and downstream performance: less feature collapse leads to more transferable models. However, it remains unclear why features prior to projection heads exhibit less collapse and greater transferability. Our study addresses this question and offers a theoretical insight into the utilization of projection heads. According to Theorem 4, it becomes apparent that features from the final layers tend to be more collapsed than the features before the projection head. This, together with the empirical correlation between feature collapse and transferability, implies that leveraging features before projection heads improves transferability. We provide further empirical evidence with experiments in Section 5.3.2. Moreover, the progressive compression pattern in Theorem 4 also provides insight into the phenomenon studied in yosinski2014transferable, which suggests that deeper layers in a neural network become excessively specialized for the pre-training task, consequently limiting their effectiveness in transfer learning.
3.2 Discussions on Assumption 3
In this subsection, we justify the properties of trained network weights in Assumption 3 using both theoretical and experimental findings. While the solution described in Assumption 3 is only one of infinitely many global optima for Problem (3), the implicit bias of GD ensures that, with proper initialization, iterations almost always converge to the desired solution. In the following discussion, we delve into this phenomenon in greater depth.
@@ 3.2 Discussions on Assumption 3
$\bm{W}_{l}{(0)^{T}}\bm{W}_{l}{(0)}=\xi^{2}\bm{I}_{d},\forall l\in[L-1],\quad\bm{W}_{L}{(0)}=\left[\xi\bm{U}\ \bm{0}\right],$ (18)
where $\bm{U}\in\mathcal{O}^{K}$. It is worth noting that orthogonal initialization is widely used to train deep networks, which can speed up the convergence of GD; see, e.g., pennington2018emergence; xiao2018dynamical; Hu2020.
Theoretical justification of Assumption 3.
Theoretically, we can prove the conditions outlined in Assumption 3 when the gradient flow is trained on a square and orthogonal data matrix using the results in arora2018optimization; yaras2023law; yaras2024compressible; kwon2023compressing.
Proposition 5
@@
The proof of this proposition is deferred to Section A.5. Although we can only prove Assumption 3 in this restrictive setting, we see empirically that it seems to hold in more general settings. We substantiate this claim by drawing support from empirical evidence and existing findings in the literature, as elaborated in the following.
Empirical justifications of Assumption 3.
@@
Figure 7: Assumption 3 holds approximately among different network configurations. We train DLNs using the orthogonal initialization with varying widths and depths. For each network width $d$ and depth $L$, we show the minimum-norm residual, the average weight balancedness residual of all weights, and the variance of the singular values of all weights (see Section 5.2.1 for details) for the DLNs that have been trained until convergence. Each data point represents the mean value of 10 runs with different random seeds, with error bars indicating standard deviation. Here, we plot these metrics in (a) with fixed network width $d$ and varying depth $L$ and (b) with fixed network depth $L=5$ and varying network width $d$. More experimental details are deferred to Section 5.2.1.
Prior arts on implicit bias support Assumption 3.
Moreover, Assumption 3 is also well supported by many related results in the literature.
  • Minimum-norm solutions and balanced weights. It has been extensively studied in the literature that GD with proper initialization converges to minimum-norm and balanced solutions, especially in the setting of gradient flow (arora2018optimization; chatterji2023deep; du2018algorithmic; min2021explicit). Specifically, du2018algorithmic; arora2018optimization proved that for linear networks, the iterates of gradient flow satisfy
@@
    where $t$ denotes a continuous time index. This, together with (18), implies that (11) holds with $\delta=\xi\sqrt{d-K}$. Moreover, using the result in min2021explicit, if we initialize the weights of two-layer networks as in (18) with $\bm{W}_{2}(0)\bm{W}_{1}(0)\in\mathrm{span}(\bm{X})$, gradient flow always yields the minimum-norm solution upon convergence. However, the conditions (10) and (11) are rarely studied in the context of GD for training deep networks. In this context, we empirically verify these two conditions as shown in Figure 7 under different network configurations.
  • Approximate low-rankness of the weights. Recently, numerous studies have demonstrated that GD exhibits a bias towards low-rank weights (arora2019implicit; Gunasekar2017; yaras2023law; kwon2023compressing). In particular, yaras2023law; kwon2023compressing showed that when the input data $\bm{X}$ is orthonormal, learning dynamics of GD for DLNs in (17), with the initialization in (18), only updates an invariant subspace of dimension $2K$ for each weight matrix across all layers, where the invariant subspace is spanned by the singular vectors associated with the $K$-largest and $K$-smallest singular values, and the rest singular values remain unchanged. More experimental demonstration can be found in Appendix B.
@@ 4 Relationship to Prior Arts
Deep networks, organized in hierarchical layers, can perform effective and automatic feature learning (allen2023backward), where the layers learn useful representations of the data. To better understand hierarchical feature learning, plenty of studies have been conducted to investigate the structures of features learned at intermediate layers. One line of these works is to investigate the neural collapse (NC) properties at intermediate layers. For example, tirer2022extended extended the study of neural collapse to three-layer nonlinear networks with the MSE loss, showing that the features of each layer exhibit neural collapse. dang2023neural generalized the study of neural collapse for DLNs with imbalanced training data, and they drew a similar conclusion to tirer2022extended that the features of each layer are collapsed. However, these results are based on the unconstrained features model and thus cannot capture the input-output relationship of the network. Moreover, the conclusion they draw on the collapse of intermediate features is far from what we observe in practice. Indeed, it is seen for both linear and nonlinear networks in Figure 5 that the intermediate features are progressively compressed across layers rather than exactly collapsing to their means for each layer. In comparison, under the assumption that the input $\bm{X}$ is nearly orthogonal in Assumption 2, our result characterizes the progressive compression and discrimination phenomenon to capture the training phenomena on practical networks as demonstrated in Figure 5.
On the other hand, another line of work (yu2020learning; chan2022redunet; ma2022principles; yu2023white) argues that deep networks prevent within-class feature compression while promoting between-class feature discrimination, contrasting with our study where features are compressed across layers. Such a difference can be attributed to the choice of loss function. In our work, we focused on the study of the commonly used MSE loss, with the goal of understanding a prevalent phenomenon in classical training of deep networks. It has been empirically shown that increasing feature compression is beneficial for improving in distribution generalization and robustness (papyan2020prevalence; li2022principled; ben2022nearest; he2023law; Chen2022a). In comparison, yu2020learning introduced a new maximum coding rate reduction loss that is intentionally designed to prevent feature compression.
@@
Our main result in Theorem 4 is based on Assumption 3 for the trained weights. As discussed in Section 3.2, Assumption 3 holds as a consequence of results in recent works on analyzing the learning dynamics and implicit bias of gradient flow or GD in training deep networks. We briefly review the related results as follows. For training DLNs, arora2018convergence established linear convergence of GD based upon whitened data and with a similar setup to ours. For Gaussian initialization, Du2019 showed that GD also converges globally at a linear rate when the width of hidden layers is larger than the depth, whereas Hu2020 demonstrated the advantage of orthogonal initialization over random initialization by showing that linear convergence of GD with orthogonal initialization is independent of the depth. More recent developments can be found in gidel2019implicit; Nguegnang2021; shin2022effects for studying GD dynamics. On the other hand, another line of works focused on studying gradient flow for learning DLNs due to its simplicity (bah2022learning; Eftekhari2020; min2021explicit; Tarmoun2021; bah2022learning), by analyzing its convergence behavior.
Numerous studies have shown that the effectiveness of deep learning is partially due to the implicit bias of its learning dynamics, which favors some particular solutions that generalize exceptionally well without overfitting in the over-parameterized setting (Belkin2019; huh2021low; neyshabur2017implicit). To gain insight into the implicit bias of GD for training deep networks, a line of recent work has shown that GD tends to learn simple functions (cao2023implicit; Gunasekar2017; Ji2018; kunin2022asymmetric; Shah2020; VallePerez2018). For instance, some studies have shown that GD is biased towards max-margin solutions in linear networks trained for binary classification via separable data (soudry2018implicit; kunin2022asymmetric). In addition to the simplicity bias, another line of work showed that deep networks trained by GD exhibit a bias towards low-rank solutions (Gunasekar2017; huh2021low; yaras2023law). The works (gidel2019implicit; arora2019implicit) demonstrated that adding depth to matrix factorization enhances an implicit tendency towards low-rank solutions, leading to more accurate recovery.
@@ 5 Experimental Results
In this section, we conduct various numerical experiments to validate our assumptions, verify our theoretical results, and investigate the implications of our results on both synthetic and real data sets. All of our experiments are conducted on a PC with 8GB memory and an Intel(R) Core i5 1.4GHz CPU, except for those involving large datasets such as CIFAR and FashionMNIST, which are conducted on a server equipped with NVIDIA A40 GPUs. Our code is implemented in Python and made available at https://github.com/Heimine/PNC_DLN. Throughout this section, we will repeatedly use multilayer perception (MLP) networks, where each layer of MLP networks consists of a linear layer and a batch norm layer (ioffe2015batch) followed by ReLU activation. In addition, in accordance with our theoretical analysis, the compression metric is plotted for layers $0,1,\dots,L-1$, while the discrimination metric is plotted for layers $1,\dots,L-1$.
The remaining sections are organized as follows. In Section 5.1, we provide detailed experimental setups for the results discussed in Section 1 and shown in Figures 6, 8, and 9. In Section 5.2, we provide experimental results to support Theorem 4 in Section 3 and validate Assumption 3 discussed in Section 3.2. Finally, we empirically explore the implications of our results in Section 5.3 beyond our assumptions.
@@ 5.1 Significance of Studying DLNs in Feature Learning
(a) A 8-layer MLP network with ReLU
(b) A 8-layer hybrid network
Figure 8: Accuracy and numerical rank of two 8-layer networks before training. We use the same setup as in Figures 1 and 6 but instead plot the numerical rank and training accuracy of the two networks across layers before training.
(a) A 8-layer MLP network with ReLU
(b) A 8-layer hybrid network
Figure 9: Within-class compression and Between-class discrimination of two 8-layer networks before training. We use the same setup as in Figure 1 and Figure 6 but instead plot the compression and discrimination statistics of the two networks across layers before training.
In this subsection, our goal is to empirically demonstrate the significance of studying DLNs for feature learning as discussed in Section 1. To achieve this, we investigate (i) the roles of linear layers and MLP layers at deep layers in feature learning and (ii) the importance of the depth of DLNs for generalization.